OpenStudy (anonymous):

On a 56 question exam, Leslie answered 7 questions incorrectly. What percent of the questions did she answer correctly. (I don't understand how the answer is 87.5%..I thought it'd be 88% !

OpenStudy (anonymous):

56-7=49 49/56 = .875 = 87.5%

OpenStudy (anonymous):

So Leslie got 49 right out of 56, then 49 / 56 = 87.5%

OpenStudy (anonymous):

Ok wow, now I know what I did wrong, thank you so much!